### v1

@@ 3.1 Inconsistency of Effect Direction
The critical condition within this proposition, “$\exists j$ and $k$, such that $r^{j}_{k}(x_{1})\neq r^{j}_{k}(\tilde{x}_{1})$,” elucidates a scenario where altering the treatment level affects the salience of at least one attribute within one profile. In the accompanying proof (detailed further in the SI 2), it is demonstrated that such a reversal can occur irrespective of the utility magnitudes associated with excluded attributes, or the inherent salience of the treatment attribute ($X_{1}$). This finding has profound implications for experimental design and interpretation, particularly in contexts where the primary interest lies in understanding the direction of causal effects. It warns researchers of the possibility that, under the influence of salience effects, the observed direction may actually invert.
@@
Based on Proposition 2, we formulate the following testable hypothesis:
@@
The direction of attribute effects reverses with the different numbers of attributes in a conjoint experiment.
We do not assert that an effect sign reversal will inevitably occur. For instance, if one attribute—or its salience—predominantly overshadows others in the evaluation process, the reversal of this attribute’s effect is unlikely to be observed. The reasoning behind this is straightforward: if a particular attribute (consider $X_{1}$) is significantly more important than any other attributes, it will be assigned a very large weight $\alpha_{1}$, and, typically, it will also possess a large utility value $u(X_{1})$. Therefore, in the evaluation $V$, even if the order of salience changes, the product $\alpha_{1}u(X_{1})$ remains considerably large, overshadowing other terms in the evaluation. The direct consequence is that its sign will predominantly determine the sign of $V$, and thus the ITE.
Another scenario that could preclude the observation of an effect sign reversal is the non-fulfillment of the rank change assumption. As highlighted by Proposition 2, the condition of salience rank change is critical for the observed result. The proposition below clarifies that without this condition, effect sign reversal cannot be observed under any circumstances.
@@
Attribute effect direction reversal is less likely if the level of attributes changes only marginally.
3.1.1 Testing the Effect Reversal Hypotheses
@@
Figure 4 presents a heatmap of the results (Figure A.3 in SI labels statistical significance). The horizontal axis represents the number of attributes, and the vertical axis denotes the level of each attribute. Dark colors indicate negative AMCEs, while light colors signify positive AMCEs. The results corroborate our hypothesis on effect sign reversal (Hypothesis 2). For instance, the effects of certain variables, such as the menu, bar, closet, or pillow, are notably unstable. However, others, including View, Towels, and Internet, demonstrate considerable stability. This observation aligns with our theoretical framework, which posits that these variables are among the most critical factors in decision-making evaluations of a hotel, possessing sufficiently high prior salience and utility to dominate over other attributes. This evidence further solidifies the notion that View and Internet are indeed core attributes, as described by Bansak et al. (2021).
@@ 3.2.1 Testing the hypothesis
To test the hypothesis, we utilize the previously mentioned data from the hotel rooms conjoint experiments. Recall that there are 11 experiment groups, featuring 4, 5, 6, 7, 8, 9, 10, 12, 14, 18, or 22 attributes. For each group, we estimate the AMCE and corresponding p values. To facilitate a clearer visualization of the results, we assign each attribute a rank based on its absolute magnitude within each group. In the heatmap 5 (and (figure A.4 in SI), each column presents the relative importance rank of attributes in a specific experiment group, as listed on the x-axis. The colors denote the importance rank. Also, we construct importance by $R^{2}$ as Bansak et al. (2018). Figure A.5 in SI shows the similar results.
@@ 4 Discussion
Firstly, it is essential to acknowledge that salience effects are not inherently detrimental to experimental design. Whether individuals are making decisions in the real world or within an experimental framework, salience effects invariably play a role in multi-dimensional decision-making processes. As such, the objective should not be to entirely eliminate salience effects from experiments. Rather, our focus should be on ensuring that the salience effects present in the experiment closely mirrors those encountered in real-world scenarios. It is the artificially induced salience effects that pose problems. Note that in the absence of attention effects—implying that the experiment perfectly replicates the real-world decision-making environment—experimental ITEs would align with real-world effects due to the accurate matching of salience effects. However, the main challenge is that achieving this ideal scenario is often impractical. Furthermore, our focus on Average Effects (AMCE) rather than Individual Effects (ITE) necessitates a particular consideration. As outlined in the SI 4, AMCEs are derived by marginalizing over attribute components and individuals, typically requiring participants to make several rounds of decisions among multiple profiles in a conjoint experiment, with each round featuring randomly generated profiles. If the randomly assigned attribute levels are uniquely distinct and lack counterparts in the real world, this can lead to distorted salience effects. This observation leads to our first recommendation for experimental design.
@@
Secondly, in our model, attention effects are a primary cause of the inconsistencies observed in experimental outcomes. Our propositions are fundamentally anchored in the assumption of limited attention, which posits that the attributes included in an experiment may not align with those deemed relevant by DMs in real-world settings. An immediate implication for research design is that researchers should endeavor to identify which attributes DMs truly value in real-life situations, thereby ensuring a match between the consideration sets in experiment and real-world contexts. As shown by Hainmueller, Hangartner and Yamamoto (2014), if the consideration sets are matched, survey experiments have ecological validity. This alignment requires a meticulous preliminary survey prior to the experimental design. It is important to note that this does not advocate for the inclusion of an excessive number of attributes, especially those that are superfluous. While the presence of an unnecessary attribute may be harmless if it has no impact on the evaluation function, even a minimal effect from such an attribute can significantly distort outcomes when coupled with salience effects. For example, when a DM encounters an unexpected attribute in a conjoint experiment, this surprise can lead the DM to assign disproportionately high salience to this attribute, despite its negligible importance in a real-world context. This observation leads us to our second recommendation for experimental design:
@@
Pilot testing of the experiment with a select sample of participants from the target population would serve as a valuable step in this process. Feedback obtained from the pilot test could be used for fine-tuning the relevance and realism of the attributes and profiles. Based on this feedback, necessary adjustments can be made to the experimental design to ensure its close resemblance to real-world decision-making contexts.
In the final stage of the experimental design process, the feedback from the pilot test and expert validation is integrated to refine the design. The culmination of this process is an experimental design that is not only aligned with real-world decision-making contexts but also equipped to provide meaningful insights into the dynamics of real-world decision-making processes.
@@ 5 Concluding Remarks
An increasing number of studies employ experiments to establish causal effects within the social sciences (Druckman et al. 2006). Beyond the assurance of internal validity, external validity remains a longstanding concern. This article delves into a more fundamental aspect of generalizability: ecological validity. We theoretically and empirically demonstrate that experiments engaging in multi-dimensional decision-making processes, especially conjoint survey experiments, may lack ecological validity. In particular, we find that the magnitude of experimental effects can be inflated, and the direction of effects and their relative importance may diverge from those observed in real-world scenarios.
Through the lens of formal theory, this study uncovers how attention and salience effects could undermine ecological validity. The alignment between experimental conditions and real-world scenarios lays the groundwork for an in-depth exploration of external validity, involving variations in treatment, outcome measures, populations, and contexts. Developing a more comprehensive theory that integrates ecological validity with other forms of external validity represents a promising direction for future research.
@@ Appendix D Design of Candidate-Choice Experiment
This study is pre-registered with the center for Open Science Framework (OSF). An anonymized version along with the questionnaire can be found at the OSF). The survey experiment received full Institutional Review Board (IRB) approval in Study ID [details removed for anonymous peer review] from the authors’ institutions.
@@ D.2 Pre-registered Design
The online anonymized version along with the questionnaire can be found at the OSF).

### v2

@@ 3.1 Inconsistency of Effect Direction
The critical condition within this proposition, “$\exists j$ and $k$, such that $r^{j}_{k}(x_{1})\neq r^{j}_{k}(\tilde{x}_{1})$,” elucidates a scenario where altering the treatment level affects the salience of at least one attribute within one profile. In the proof (detailed further in the SI 2), it is demonstrated that such a reversal can occur irrespective of the utility magnitudes associated with excluded attributes, or the inherent salience of the treatment attribute ($X_{1}$). This finding has profound implications for experimental design and interpretation, particularly in contexts where the primary interest lies in understanding the direction of causal effects.
@@
A weaker conclusion implied by this example and proposition 2 is that the effect observed in one scenario (the experimental world) may not necessarily reverse in a different scenario (the real world); instead, it may simply dissipate to no effect. Intuitively, if parameters exist so that they can support the case that pulls down a positive effect to be negative, it is even more likely that parameters could be found to support the case that reducing the effect to null. We formalize this observation in the following corollary:
Corollary 2.
Suppose assumptions 2 and 3 hold. $\exists j$ and $k$, such that $r^{j}_{k}(x_{1})\neq r^{j}_{k}(\tilde{x}_{1})$, then experimental treatment effect may become a null effect in the real world, and vice versa.
Based on Proposition 2 and corollary 2, we formulate the following testable hypothesis:
@@
The sign of the effect in a conjoint experiment may reverse or diminish to a null effect depending on the number of attributes involved.
We do not assert that an effect sign reversal or diminish will inevitably occur. For instance, if one attribute—or its salience—predominantly overshadows others in the evaluation process, the reversal of this attribute’s effect is unlikely to be observed. The reasoning behind this is straightforward: if a particular attribute (consider $X_{1}$) is significantly more important than any other attributes, it will be assigned a very large weight $\alpha_{1}$, and, typically, it will also possess a large utility value $u(X_{1})$. Therefore, in the evaluation $V$, even if the order of salience changes, the product $\alpha_{1}u(X_{1})$ remains considerably large, overshadowing other terms in the evaluation. The direct consequence is that its sign will predominantly determine the sign of $V$, and thus the ITE.
Another scenario that could preclude the observation of an effect sign reversal or diminish is the non-fulfillment of the rank change assumption. As highlighted by Proposition 2, the condition of salience rank change is critical for the observed result. The proposition below clarifies that without this condition, effect sign reversal cannot be observed under any circumstances.
@@
Attribute effect direction reversal or diminish is less likely if the level of attributes changes only marginally.
3.1.1 Test Hypotheses 2 and 3
@@
Figure 3 presents a heatmap of the results [^9] Figure A.3 in SI labels statistical significance. Due to low statistical power, we observe a limited number of significant estimates. As the sample size increases, we anticipate that the confidence intervals will narrow around the estimates, potentially leading to observations of significant effect reversals. However, observing a null effect is also consistent with our theoretical results and aligns with Hypothesis 2.. The horizontal axis represents the number of attributes, and the vertical axis denotes the level of each attribute. Dark colors indicate negative AMCEs, while light colors signify positive AMCEs. The results corroborate our hypothesis on effect sign reversal (Hypothesis 2). For instance, the effects of certain variables, such as the menu, bar, closet, or pillow, are notably unstable. However, others, including View, Towels, and Internet, demonstrate considerable stability. This observation aligns with our theoretical framework, which posits that these variables are among the most critical factors in decision-making evaluations of a hotel, possessing sufficiently high prior salience and utility to dominate over other attributes. This evidence further solidifies the notion that View and Internet are indeed core attributes, as described by Bansak et al. (2021).
@@ 3.2.1 Testing the hypothesis
To test the hypothesis, we utilize the previously mentioned data from the hotel rooms conjoint experiments. Recall that there are 11 experiment groups, featuring 4, 5, 6, 7, 8, 9, 10, 12, 14, 18, or 22 attributes. For each group, we estimate the AMCE and corresponding p values. To facilitate a clearer visualization of the results, we assign each attribute a rank based on its absolute magnitude within each group. In the heatmap 4, each column presents the relative importance rank of attributes in a specific experiment group, as listed on the x-axis. The colors denote the importance rank. [^10] Figure A.4 in SI indicates statistical significance. Also, we construct importance by $R^{2}$ as Bansak et al. (2018). Figure A.5 in SI shows the similar results.
@@ 4 Discussion
Firstly, it is essential to acknowledge that salience effects are not inherently detrimental to experimental design. Whether individuals are making decisions in the real world or within an experimental framework, salience effects invariably play a role in multi-dimensional decision-making processes. As such, the objective should not be to entirely eliminate salience effects from experiments. Rather, our focus should be on ensuring that the salience effects present in the experiment closely mirrors those encountered in real-world scenarios. It is the artificially induced salience effects that pose problems. Note that in the absence of attention effects—implying that the experiment perfectly replicates the real-world decision-making environment—experimental ITEs would align with real-world effects due to the accurate matching of salience effects. However, the main challenge is that achieving this ideal scenario is often impractical. Furthermore, our focus on Average Effects (AMCE) rather than Individual Effects (ITE) necessitates a particular consideration. As outlined in the SI 3, AMCEs are derived by marginalizing over attribute components and individuals, typically requiring participants to make several rounds of decisions among multiple profiles in a conjoint experiment, with each round featuring randomly generated profiles. If the randomly assigned attribute levels are uniquely distinct and lack counterparts in the real world, this can lead to distorted salience effects. This observation leads to our first recommendation for experimental design.
@@
Secondly, in our model, attention effects are a primary cause of the inconsistencies observed in experimental outcomes. Our propositions are fundamentally anchored in the assumption of limited attention, which posits that the attributes included in an experiment may not align with those deemed relevant by DMs in real-world settings. An immediate implication for research design is that researchers should endeavor to identify which attributes DMs truly value in real-life situations, thereby ensuring a match between the consideration sets in experiment and real-world contexts. As shown by Hainmueller, Hangartner and Yamamoto (2014), if the consideration sets are matched, survey experiments have consistency. This alignment requires a meticulous preliminary survey prior to the experimental design. It is important to note that this does not advocate for the inclusion of an excessive number of attributes, especially those that are superfluous. While the presence of an unnecessary attribute may be harmless if it has no impact on the evaluation function, even a minimal effect from such an attribute can significantly distort outcomes when coupled with salience effects. For example, when a DM encounters an unexpected attribute in a conjoint experiment, this surprise can lead the DM to assign disproportionately high salience to this attribute, despite its negligible importance in a real-world context. This observation leads us to our second recommendation for experimental design:
@@ 5 Concluding Remarks
An increasing number of studies employ experiments to establish causal effects within the social sciences (Druckman et al. 2006). Beyond the assurance of internal validity, external validity remains a longstanding concern. This article delves into a more fundamental aspect of generalizability: the consistency of experimental and real-world effects. We theoretically and empirically demonstrate that experiments engaging in multi-dimensional decision-making processes, especially conjoint survey experiments, may lack such consistency. In particular, we find that the magnitude of experimental effects can be inflated, and the direction of effects and their relative importance may diverge from those observed in real-world scenarios.
A caveat to our findings is that we do not claim the inconsistency problem is inevitable in all such experiments. Many studies, such as those by Jenke et al. (2021), along with empirical examples discussed in our paper, suggest the stable consistency of certain attributes. Instead, we propose two plausible mechanisms and outline conditions that could lead to inconsistencies. From an empirical perspective, this encourages researchers to be more vigilant when designing experiments. Theoretically, we incorporate behavioral theories to deepen our understanding of experimental dynamics. The alignment between experimental conditions and real-world scenarios lays the groundwork for an in-depth exploration of external validity, involving variations in treatment, outcome measures, populations, and contexts. Developing a more comprehensive theory that integrates consistency with other forms of external validity represents a promising direction for future research.
@@ Appendix D Design of Candidate-Choice Experiment
This study is pre-registered with the center for Open Science Framework (OSF). An anonymized version along with the questionnaire can be found at the OSF. The survey experiment received full Institutional Review Board (IRB) approval in Study ID [details removed for anonymous peer review] from the authors’ institutions.
@@ D.2 Pre-registered Design
The online anonymized version along with the questionnaire can be found at the OSF.